OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

slope of the line passing through those two pts is -1/5. (slope = y2 - y1 / x2 - x1 ) now plug in the values in the slope-point form of equation: y - y1 = m ( x - x1) u know the value of m, take one of the given pts as (x1, y1).